What is the answeror solution of x+2y=6

Answer by macston(5194)

.
You can not solve for two variables with one equation.
What do you want to solve for?
.
Solve for x:
x+2y=6
x=6-2y
.
Solve for y:
x+2y=6
2y=6-x
y=(6-x)/2